Jeremy is 3 years older than Emily. In 6 years the sum of their ages will be 75. How old is Jeremy now?

Answer by FrankM(1040)

J=E+3
(J+6)+(E+6)=75
(E+3+6)+(E+6)=75
2E+15=75
E= 30
J= 33